# Ranking Entities along Conceptual Space Dimensions with LLMs: An Analysis of Fine-Tuning Strategies

### v1

@@ Abstract
Conceptual spaces represent entities in terms of their primitive semantic features. Such representations are highly valuable but they are notoriously difficult to learn, especially when it comes to modelling perceptual and subjective features. Distilling conceptual spaces from Large Language Models (LLMs) has recently emerged as a promising strategy. However, existing work has been limited to probing pre-trained LLMs using relatively simple zero-shot strategies. We focus in particular on the task of ranking entities according to a given conceptual space dimension. Unfortunately, we cannot directly fine-tune LLMs on this task, because ground truth rankings for conceptual space dimensions are rare. We therefore use more readily available features as training data and analyse whether the ranking capabilities of the resulting models transfer to perceptual and subjective features. We find that this is indeed the case, to some extent, but having perceptual and subjective features in the training data seems essential for achieving the best results. We furthermore find that pointwise ranking strategies are competitive against pairwise approaches, in defiance of common wisdom.
@@ 1 Introduction
  • We compare pointwise and pairwise approaches for ranking entities with LLMs. Despite the fact that pairwise approaches have consistently been found superior for LLM-based document ranking Nogueira et al. (2019); Gienapp et al. (2022); Qin et al. (2023), when it comes to ranking entities, we find the pointwise approach to be highly effective.
@@ 2 Related Work
Our focus in this paper is on extracting knowledge from language models. This idea of language models as knowledge bases was popularised by Petroni et al. (2019), who showed that the pre-trained BERT model captures various forms of factual knowledge, which can moreover be extracted using a simple prompt. Work in this area has focused on two rather distinct goals. On the one hand, probing tasks, such as the one proposed by Petroni et al. (2019), have been used as a mechanism for analysing and comparing different language models. On the other hand, extracting knowledge from LMs has also been studied as a practical tool for building or extending symbolic knowledge bases. This has been particularly popular for capturing types of knowledge which are not commonly found in traditional knowledge bases, such as commonsense knowledge Bosselut et al. (2019); West et al. (2022); Yu et al. (2023). Several works have focused on distilling KGs from language models Cohen et al. (2023). Hao et al. (2023) studies this problem for non-traditional relations such as “is capable of but not good at”. Along the same lines, Ushio et al. (2023) have focused on modelling relations that are a matter of degree, such as “is a competitor of” or “is similar to”. We can similarly think of the conceptual space dimensions that we consider in this paper as gradual properties.
@@
There is an ongoing debate about the extent to which LMs can truly capture meaning Bender and Koller (2020); Abdou et al. (2021); Patel and Pavlick (2022); Søgaard (2023). Within this context, several authors have analysed the ability of LMs to predict perceptual features. As already mentioned, Paik et al. (2021) and Liu et al. (2022a) analysed the ability of LMs to predict colour terms. Abdou et al. (2021) analyses whether the representation of colour terms in LMs can be aligned with their representation in the CIELAB colour space. Patel and Pavlick (2022) similarly showed that LLMs can generate colour terms from RGB codes in a few-shot setting, even if the codes represent a rotation of the standard RGB space. They also show a similar result for terms describing spatial relations. Zhu et al. (2024) have similarly shown that LLMs can understand colour codes, by using them to generate HSL codes for everyday objects, or by asking models to choose the most suitable code among two alternatives.
Beyond the colour domain, Li et al. (2023) considered physical properties such as height or mass. While they found LLMs to struggle with such properties, Chatterjee et al. (2023) reported better results on the same datasets, especially for GPT-4. Focusing on visual features, Merullo et al. (2023) showed that the representations of concepts in vision-only and text-only models can be aligned using a linear mapping. Chatterjee et al. (2023) focused on the taste domain, modelling properties such as sweet. They found that GPT-3 can model such properties to a reasonable extent, but not better than a fine-tuned BERT model.
Gupta et al. (2015) already considered the problem of modelling gradual properties in the context of static word embeddings, although their analysis was limited to objective numerical features. Derrac and Schockaert (2015) similarly learned conceptual space dimensions for properties such as “violent” in a semantic space of movies. These approaches essentially learn a linear classifier or regression model for each property indepdently, and can thus not generalise to new properties.
@@ 3.1 Pairwise Model
We typically want to rank a given set of entities, rather than judging the relative position of two particular elements. This means that we need a strategy for aggregating (noisy) pairwise judgments into a single ranking. This problem has received extensive attention in the literature, with standard techniques including spectral ranking Vigna (2016) and maximum likelihood estimation w.r.t. an underlying statistical model. However, existing approaches often consider a stochastic setting, where we may have access to several judgments for the same entity pair (e.g. when ranking sports teams based on the outcomes of head-to-head matches).
@@ 3.2 Pointwise Model
where $t_{ij}=1$ if $f(e_{i})>f(e_{j})$ and $t_{ij}=0$ otherwise, and the summation ranges over all distinct entity pairs $e_{i},e_{j}$ within the given mini-batch. Note that while we use pairwise comparisons for training the model, it is still a pointwise approach as it produces scores for individual entities.
@@ 3.3 Baselines
To put the performance of the fine-tuning strategies from Sections 3.1 and 3.2 into context, we compare them with two conversational models: ChatGPT (gpt-3.5-turbo) and GPT-4 (gpt-4). We use both models in a zero-shot setting. For this purpose, we use the same prompt as in Section 3.1 but append the sentence Only answer with yes or no. Despite this instruction, the models occasionally still generates a different response, typically expressing that the question cannot be answered. For such entity pairs, we replace the generated response with a randomly generated label (yes or no). [^1] Statistics about how often this was needed can be found in the appendix.
@@ 4 Datasets
We have obtained 20 rankings from numerical features that are available on Wikidata [^2] https://www.wikidata.org/wiki/Wikidata:Main_Page. For instance, we obtained a ranking of rivers by comparing their length. [^3] The entity types and corresponding features are listed in the appendix. If there were more than 1000 entities with a given feature value, we selected the most 1000 popular entities. To estimate the popularity of an entity, we use their QRank [^4] https://qrank.wmcloud.org, which counts the number page views of the corresponding entry in sources such as Wikipedia. For the entity type person, we limited the analysis to people born in London (which made it possible to retrieve the required information from Wikidata more efficiently). We similarly only considered museums located in Italy. For some experiments, we split the collected data in two datasets, called WD1 and WD2. This will allow us to test whether models trained on one set of features (i.e. WD1) generalise to a different set of feature (i.e. WD2). WD1 contains rankings which were cut off at 1000 elements, whereas WD2 contains rankings with fewer elements. We will write WD to refer to the full dataset, i.e. WD1 and WD2 combined.
@@
WD+TG+Rocks - - 74.0 72.4 60.0 60.2 70.6 72.2 - - - - - - - - - 85.6 88.2 62.0
WD+Taste+Rocks - - - - - - - - - - - - - - - 69.1 65.2 89.4 91.2 63.8
@@
Taste
Rocks
Tag Genome
Vig et al. (2012) collected a dataset [^6] Available from https://grouplens.org/datasets/movielens/tag-genome-2021/. of movies, called the Tag Genome, by asking annotators to what extent different tags apply to different movies, on a scale from 1 to 5. From these tags, we first selected those that correspond to adjectives and for which ratings for at least 15 movies were available. We then manually identified 38 of these adjectives which correspond to ordinal features. More recently, Kotkov et al. (2022) created a similar dataset for books. We again selected adjectives for which at least 15 items were ranked, and manually identified 32 adjectives that correspond to ordinal features. A list of the adjectives that we considered, together with the corresponding number of items in the rankings, is provided in the appendix. It should be noted that most items are only judged by a single annotator, and the judgements were moreover obtained using crowdsourcing. The movies and books datasets are thus clearly noisier than the taste and rocks datasets. By averaging across a large number of rankings, we believe that these datasets can nonetheless be valuable. For this reason, we will only consider aggregated results across all tags when evaluating on these datasets. We will write TG to refer to the combined dataset, containing both the books and movies rankings.
Physical Properties
@@
Pointwise 51.0 64.8 32.5 35.2 52.0 61.7
SVM (5 samples) 62.1 62.2 42.6 44.6 56.4 63.0
SVM (30 samples) 66.0 64.7 47.6 47.0 60.6 65.8
Count (5 samples) 59.0 57.4 46.7 41.7 53.5 60.1
Count (30 samples) 64.8 64.7 49.1 47.2 59.9 64.7
$*{}^{*}$ 17.5 8.5 12.2 16.4 22.5 10.7
$*{}^{*}$ 19.5 51.1 20.2 22.0 22.6 16.0
$*{}^{*}$ 36.0 46.3 32.8 23.2 22.6 31.7
$*{}^{*}$ 55.0 63.2 33.3 27.2 57.0 52.0
Table 3: Comparison of ranking strategies, in terms of Spearman $\rho$%. Baseline results marked with $*{}^{*}$ were taken from Chatterjee et al. (2023). All other results are obtained with Llama-13B trained on WD+TG+Rocks.
@@
Sweetness
mango, dried date, white chocolate, peach , pineapple in syrup, fruit candy, syrup with water, ice cream, strawberry, sweet pancake with maple syrup
minced beef patty, grilled calf livers, squid, sandwich with cold cuts, gizzards, croque-monsieur, roast rabbit, stir-fried bacon, roast beef , calf head with vinaigrette
Saltiness
green olives, extruded salty crackers, soy sprouts with soy sauce, canned anchovies, canned sardines, pasta with soy sauce, salted pies, marinated mussels, potato chips, salted cake
clafoutis, raspberry cake, stewed apple, raspberry with whipped cream, white chocolate, strawberry with cream and sugar, mix fruits juice, apple, raspberry, strawberry
Scary
Descent, The (2005), Grudge, The (2004), Exorcist, The (1973), Silence of the Lambs, The (1991), Ring, The (2002), Texas Chainsaw Massacre, The (1974), Shining, The (1980), Seven (a.k.a. Se7en) (1995), Amityville Horror, The (2005), American Werewolf in London, An (1981)
Super Size Me (2004), Station Agent, The (2003), Ray (2004), Dances with Wolves (1990), Jerry Maguire (1996), Driving Miss Daisy (1989), School of Rock (2003), Kung Fu Panda (2008), Miss Congeniality (2000), Ninotchka (1939)
Funny
Ace Ventura: When Nature Calls (1995), Ace Ventura: Pet Detective (1994), Hot Shots! Part Deux (1993), Army of Darkness (1993), South Park: Bigger, Longer and Uncut (1999), Auntie Mame (1958), Blazing Saddles (1974), Clerks (1994), Grand Day Out with Wallace and Gromit, A (1989), Hitchhiker’s Guide to the Galaxy, The (2005)
Spanish Prisoner, The (1997), Son of Dracula (1943), Ghost Dog: The Way of the Samurai (1999), Ferngully: The Last Rainforest (1992), High Crimes (2002), Cadillac Man (1990), Bad Boys II (2003), House of Wax (1953), Fire in the Sky (1993), Step Up 2 the Streets (2008)
Population
India, Nigeria, People’s Republic of China, Iran, Pakistan, United States of America, Russia, Indonesia, Egypt, Bangladesh
Dominica, Nauru, Andorra, Cook Islands, Saint Vincent and the Grenadines, Seychelles, Palau, Northern Mariana Islands, Liechtenstein, Niue
Table 4: We show the top and bottom ranked entities for five features: sweetness and saltiness, from the food dataset, scary and funny, from movies, and countries population, from WD2.
@@ 5 Experiments
Table 1 compares a number of different models. We test three different LLMs: the 7B and 13B parameter Llama 2 models [^8] We use the llama-2-7b-hf and llama-2-13b-hf models available from https://huggingface.co/meta-llama. and the 7B parameter Mistral model [^9] We use the mistral-7b-v0.1 model available from https://huggingface.co/mistralai/Mistral-7B-v0.1.. We evaluate the different models in terms of their accuracy on pairwise judgements. To this end, for a given dataset, we randomly sample pairs of entities $e_{i},e_{j}$ and construct queries asking whether $f(e_{i})>f(e_{j})$. For WD, Taste and Rocks, we sample 500 such pairs for each of the features. Since the TG dataset has a total of 70 features, we limit the test set to 100 pairs per feature. For this analysis, we have split the WD1 dataset into two parts: 80% of the entities, for each feature, are used for training the models. The remaining 20% are used as a test set. All models are fine-tuned on the training split of WD1 (apart from the baselines, which are evaluated zero-shot). This allows us to see how well the models perform on the features they were trained on (by evaluating on the WD1 test split), as well as how they generalise to unseen properties.
The aim of the analysis in Table 1 is to assess whether models can be successfully fine-tuned using a relatively small training set (i.e. WD1-train), involving only well-defined numerical features. In particular, we want to test whether models which are fine-tuned on such features would also generalise to more subjective and less readily available ones, similar to the easy-to-hard generalisation that has been observed for LLMs in other tasks Hase et al. (2024). The results show that this is only the case to some extent. Overall, we can see that Mistral-7B achieves the best results among the pointwise models, while Llama2-13B achieves the best results among the pairwise models. The performance of the pointwise Mistral-7B model is particularly surprising, given that pairwise models generally perform better in ranking tasks. The performance of the models across different features is not always consistent. Each model achieves close to random chance on some of the features, but the features where one model performs poorly are not always the same features where other models perform poorly. However, for bitterness and roughness, all models perform below 60% F1. Furthermore, sourness and organisation also stand out as being more challenging. Regarding the baselines, GPT-4 generally performs better than the fine-tuned models. ChatGPT performs worse on most features, but achieves the best results for books.
@@
The relatively disappointing results from Table 1 can be partially explained by the fact that a small training set was used, which moreover only covered numerical features and particular entity types. In Table 2, we evaluate the impact of using different training sets. For this analysis, we use the pairwise Llama2-13B model. Our focus is on seeing whether models trained on one domain can generalise to other domains. The results are again evaluated in terms of accuracy, using the same pairwise judgements as for Table 1. WD refers to the full dataset (including both the training and test splits of WD-1).
@@
We can see that training on larger datasets indeed leads to considerably better results. While this is not unexpected, we can also make more striking observations in Table 2. For instance, the model that was trained on Taste alone achieves strong results on Rocks, despite the two datasets involving very different features. Similarly, the model that was only trained on TG achieves strong results for both Taste and Rocks. This suggests that the fine-tuned models are indeed capable of generalising to unseen domains. However, to achieve strong results, it appears to be important that the training data contains subjective or perceptual features. Indeed, training on TG alone overall performed poorly, compared to the other training sets. The best results in Table 2 are competitive with the GPT-4 results from Table 1. Given that the training and test sets cover disjoint domains, the results in Table 2 reflect the knowledge that is captured by the LLMs themselves, rather than knowledge that was injected during the fine-tuning process. This suggests that pre-trained LLMs capture more perceptual knowledge than it may initially appear.
@@
Table 3 compares different strategies for generating rankings. The pointwise model can be used directly for this purpose. For the pairwise model, we show results with the SVM strategy and the Count strategy. For Count, we furthermore vary the number of pairwise judgments per entity (5 or 30). For this experiment, we use the (pointwise and pairwise) Llama2-13B models that were trained on WD+TG+Rocks. We evaluate the different models by comparing the predicted rankings with the ground truth in terms of Spearman $\rho$. We can see that the pairwise approaches outperform the pointwise model in this case. This is somewhat surprising, given the strong performance of the pointwise models in Table 1. Essentially, because the ranking strategies aggregate many pairwise samples, the noisy nature of the pairwise judgments can to some extent be mitigated. The SVM method generally performs better than the Count method, especially in the case where only 5 judgments per entity are obtained. We also compare with the GPT-3 results reported by Chatterjee et al. (2023), finding that the pairwise Llama model consistently performs best.

### v2

@@ Abstract
Conceptual spaces represent entities in terms of their primitive semantic features. Such representations are highly valuable but they are notoriously difficult to learn, especially when it comes to modelling perceptual and subjective features. Distilling conceptual spaces from Large Language Models (LLMs) has recently emerged as a promising strategy, but existing work has been limited to probing pre-trained LLMs using relatively simple zero-shot strategies. We focus in particular on the task of ranking entities according to a given conceptual space dimension. Unfortunately, we cannot directly fine-tune LLMs on this task, because ground truth rankings for conceptual space dimensions are rare. We therefore use more readily available features as training data and analyse whether the ranking capabilities of the resulting models transfer to perceptual and subjective features. We find that this is indeed the case, to some extent, but having at least some perceptual and subjective features in the training data seems essential for achieving the best results. [^1] Our pre-processed datasets and code can be found at https://github.com/niteshroyal/RankingUsingLLMs.
@@ 1 Introduction
  • We compare pointwise and pairwise approaches for ranking entities with LLMs. Despite the fact that pairwise approaches have consistently been found superior for LLM-based document ranking Nogueira et al. (2019); Gienapp et al. (2022); Qin et al. (2023), when it comes to comparing entities, we find the pointwise approach to be highly effective, although the best results overall are still obtained using a pairwise model.
@@ 2 Related Work
Our focus in this paper is on extracting knowledge from language models. This idea of language models as knowledge bases was popularised by Petroni et al. (2019), who showed that the pre-trained BERT model captures various forms of factual knowledge, which can moreover be extracted using a simple prompt. Work in this area has focused on two rather distinct goals. On the one hand, probing tasks, such as the one proposed by Petroni et al. (2019), have been used as a mechanism for analysing and comparing different language models. On the other hand, extracting knowledge from LMs has also been studied as a practical tool for building or extending symbolic knowledge bases. This has been particularly popular for capturing types of knowledge which are not commonly found in traditional knowledge bases, such as commonsense knowledge Bosselut et al. (2019); West et al. (2022); Yu et al. (2023). Several works have focused on distilling KGs from language models Cohen et al. (2023). Hao et al. (2023) studied this problem for non-traditional relations such as “is capable of but not good at”. Along the same lines, Ushio et al. (2023) have focused on modelling relations that are a matter of degree, such as “is a competitor of” or “is similar to”. We can similarly think of the conceptual space dimensions that we consider in this paper as gradual properties.
@@
There is an ongoing debate about the extent to which LMs can truly capture meaning Bender and Koller (2020); Abdou et al. (2021); Patel and Pavlick (2022); Søgaard (2023). Within this context, several authors have analysed the ability of LMs to predict perceptual features. As already mentioned, Paik et al. (2021) and Liu et al. (2022a) analysed the ability of LMs to predict colour terms. Abdou et al. (2021) analysed whether the representation of colour terms in LMs can be aligned with their representation in the CIELAB colour space. Patel and Pavlick (2022) similarly showed that LLMs can generate colour terms from RGB codes in a few-shot setting, even if the codes represent a rotation of the standard RGB space. They also show a similar result for terms describing spatial relations. Zhu et al. (2024) have similarly shown that LLMs can understand colour codes, by using them to generate HSL codes for everyday objects, or by asking models to choose the most suitable code among two alternatives.
Beyond the colour domain, Li et al. (2023) considered physical properties such as height or mass. While they found LLMs to struggle with such properties, Chatterjee et al. (2023) reported better results on the same datasets, especially when using GPT-4. Focusing on visual features, Merullo et al. (2023) showed that the representations of concepts in vision-only and text-only models can be aligned using a linear mapping. Chatterjee et al. (2023) focused on the taste domain, modelling properties such as sweet. They found that GPT-3 can model such properties to a reasonable extent, but not better than a fine-tuned BERT model.
Gupta et al. (2015) already considered the problem of modelling gradual properties in the context of static word embeddings, although their analysis was limited to objective numerical features. Derrac and Schockaert (2015) similarly learned conceptual space dimensions for properties such as “violent” in a semantic space of movies, while Rubinstein et al. (2015) assessed the ability to predict (gradual) commonsense properties from word embeddings. More recently, Grand et al. (2022) also used linear projections to map word embeddings to dimensions capturing gradual feature (e.g. size). Erk and Apidianaki (2024) argue for the importance of using human ratings to learn such projections. Note that all these approaches essentially learn a linear classifier or regression model for each property independently, and can thus not generalise to new properties. A related line of work has focused on modelling the rankings associated with scalar adverbs Lorge and Pierrehumbert (2023) and adjectives Samir et al. (2021).
@@ 3.1 Pairwise Model
We typically want to rank a given set of entities, rather than judging the relative position of two particular elements. This means that we need a strategy for aggregating (noisy) pairwise judgments into a single ranking. This problem has received extensive attention in the literature, where standard techniques include spectral ranking Vigna (2016) and maximum likelihood estimation w.r.t. an underlying statistical model. However, existing approaches often consider a stochastic setting, where we may have access to several judgments for the same entity pair (e.g. when ranking sports teams based on the outcomes of head-to-head matches).
@@ 3.2 Pointwise Model
where $t_{ij}=1$ if $f(e_{i})>f(e_{j})$ and $t_{ij}=0$ otherwise, and the summation ranges over all distinct entity pairs $e_{i},e_{j}$ within the given mini-batch. Note that while we use pairwise comparisons for training the model, it is still a pointwise approach as the model produces scores for individual entities.
@@ 3.3 Baselines
To put the performance of the fine-tuning strategies from Sections 3.1 and 3.2 into context, we compare them with two conversational models: ChatGPT (gpt-3.5-turbo) and GPT-4 (gpt-4). We use both models in a zero-shot setting. For this purpose, we use the same prompt as in Section 3.1 but append the sentence Only answer with yes or no. Despite this instruction, the models occasionally still generated a different response, typically expressing that the question cannot be answered. For such entity pairs, we replace the generated response with a randomly generated label (yes or no). [^2] Statistics about how often this was needed can be found in the appendix.
@@ 4 Datasets
We have obtained 20 rankings from numerical features that are available on Wikidata [^3] https://www.wikidata.org/wiki/Wikidata:Main_Page. For instance, we obtained a ranking of rivers by comparing their length. [^4] The entity types and corresponding features are listed in the appendix. If there were more than 1000 entities with a given feature value, we selected the most 1000 popular entities. To estimate the popularity of an entity, we use their QRank [^5] https://qrank.wmcloud.org, which counts the number of page views of the corresponding entry in sources such as Wikipedia. For the entity type person, we limited the analysis to people born in London (which made it possible to retrieve the required information from Wikidata more efficiently). We similarly only considered museums located in Italy. For some experiments, we split the collected data in two datasets, called WD1 and WD2. This will allow us to test whether models trained on one set of features (i.e. WD1) generalise to a different set of features (i.e. WD2). WD1 contains rankings which were cut off at 1000 elements, whereas WD2 contains rankings with fewer elements. We will write WD to refer to the full dataset, i.e. WD1 and WD2 combined.
@@
Taste
Rocks
Tag Genome
Vig et al. (2012) collected a dataset [^7] Available from https://grouplens.org/datasets/movielens/tag-genome-2021/. of movies, called the Tag Genome, by asking annotators to what extent different tags apply to different movies, on a scale from 1 to 5. From these tags, we first selected those that correspond to adjectives and for which ratings for at least 15 movies were available. We then manually identified 38 of these adjectives which correspond to ordinal features. More recently, Kotkov et al. (2022) created a similar dataset for books. We again selected adjectives for which at least 15 items were ranked, and manually identified 32 adjectives that correspond to ordinal features. A list of the adjectives that we considered is provided in the appendix. It should be noted that most items are only judged by a single annotator, and the judgements were moreover obtained using crowdsourcing. The movies and books datasets are thus clearly noisier than the taste and rocks datasets. For this reason, we will only consider aggregated results across all tags when evaluating on these datasets. We will write TG to refer to the combined dataset, containing both the books and movies rankings.
Physical Properties
@@
Few-shot 50.8 53.6 52.2 51.8 51.2 49.2 47.8 53.4 55.2 51.4 45.2 50.8 49.2 48.6 51.4 51.5 54.9 58.8 54.2 49.6
WD1-train 83.5 66.2 68.2 79.8 61.2 56.8 72.2 69.2 58.6 75.0 60.6 53.0 54.6 58.8 73.6 62.5 56.7 94.0 94.2 63.2
WD - - 63.6 75.2 57.2 55.2 69.0 66.8 68.4 71.4 63.2 56.2 57.2 65.6 73.0 69.0 60.3 92.8 93.6 61.6
TG 68.3 68.2 75.6 80.0 62.8 58.8 71.4 69.8 70.6 73.0 58.2 62.0 58.6 58.8 73.2 - - 84.4 83.2 58.4
Taste 64.0 64.3 - - - - - - 63.2 77.2 61.6 65.0 54.0 58.4 72.6 65.2 56.0 77.4 87.8 51.2
WD+TG+Taste - - - - - - - - 69.2 80.0 64.4 59.8 59.4 65.6 73.0 - - 84.0 91.0 55.6
WD+TG+Rocks - - 74.0 81.6 66.2 57.8 73.2 70.4 - - - - - - - - - 91.2 89.4 56.8
WD+Taste+Rocks - - - - - - - - - - - - - - - 68.5 60.1 85.2 91.8 56.0
@@
Pointwise 59.3 59.9 24.0 32.1 44.1 60.4
SVM (5 samples) 60.4 65.9 48.0 38.3 59.6 62.3
SVM (30 samples) 66.4 73.5 50.9 45.4 63.9 67.2
Count (5 samples) 59.9 66.5 47.3 41.1 59.5 60.5
Count (30 samples) 65.3 72.2 48.8 44.9 63.0 66.8
Ada 17.5 8.5 12.2 16.4 22.5 10.7
Babbage 19.5 51.1 20.2 22.0 22.6 16.0
Curie 36.0 46.3 32.8 23.2 22.6 31.7
Davinci 55.0 63.2 33.3 27.2 57.0 52.0
Table 3: Comparison of ranking strategies, in terms of Spearman $\rho$%. Baseline results marked with were taken from Chatterjee et al. (2023). All other results are obtained with Llama3-8B trained on WD+TG+Rocks.
@@
Sweetness caramel custard, sweet cookies with chocolate, sweet pancake with sugar, syrup with water, macaroons, white chocolate, ice cream, litchi juice, ice cream bars, fruit candy hake in tinfoil, unsalted pasta, tripe, raw horseradish with salt, mixed salad starch based with cold cuts, gizzards, roast beef, dry sausage, pasta with salt, burbot
Saltiness dry-cured ham, parmesan cheese, canned sardines, canned anchovies, cooked ham, salty crackers, fried anchovies, cooked ham with salted butter, kidneys with gravy, gorgonzola cheese strawberry with sugar, pear, grape juice, pineapple juice, fruit crumble, strawberry juice, strawberry, banana, apple juice, blueberry
Scary Shining, The (1980), Silence of the Lambs, The (1991), Exorcist, The (1973), Psycho (1960), Ring, The (2002), Grudge, The (2004), Amityville Horror, The (2005), Texas Chainsaw Massacre, The (1974), Descent, The (2005), American Werewolf in London, An (1981) Stand by Me (1986), Kung Fu Panda (2008), Super Size Me (2004), Station Agent, The (2003), Miss Congeniality (2000), School of Rock (2003), Roger and Me (1989), Jerry Maguire (1996), Ninotchka (1939), Driving Miss Daisy (1989)
Funny Blazing Saddles (1974), This Is Spinal Tap (1984), Monty Python’s The Meaning of Life (1983), Young Frankenstein (1974), Monty Python and the Holy Grail (1975), Fish Called Wanda, A (1988), Jackass Number Two (2006), Some Like It Hot (1959), Ferris Bueller’s Day Off (1986), There’s Something About Mary (1998) Night of the Living Dead (1968), Battleship Potemkin (1925), House of Flying Daggers (Shi mian mai fu) (2004), Jesus Camp (2006), War of the Worlds (2005), Fire in the Sky (1993), Pretty Baby (1978), Ring, The (2002), Titanic (1997), United 93 (2006)
Population People’s Republic of China, India, Pakistan, Egypt, Nigeria, Russia, United States of America, Brazil, Mexico, Turkey Dominica, Niue, Tuvalu, Antigua and Barbuda, Cook Islands, Sint Maarten, Nauru, Liechtenstein, Republic of Artsakh, Andorra
Table 4: We show the top and bottom ranked entities for five features: sweetness and saltiness from the food dataset, scary and funny from movies, and countries population from WD2. Results were obtained with the pairwise Llama3-8B model trained on WD+TG+Rocks for sweetness and saltiness, trained on WD+Taste+Rocks for scary and funny, and trained on WD1 for population.
@@ 5 Experiments
Table 1 compares a number of different models. We test four different LLMs: the 7B and 13B parameter Llama 2 models, the 8B parameter Llama 3 model [^9] We use the llama-2-7b-hf, llama-2-13b-hf, and meta-llama/Meta-Llama-3-8B models available from https://huggingface.co/meta-llama., and the 7B parameter Mistral model [^10] We use the mistral-7b-v0.1 model available from https://huggingface.co/mistralai/Mistral-7B-v0.1.. We evaluate the different models in terms of their accuracy on pairwise judgements. To this end, for a given dataset, we randomly sample pairs of entities $e_{i},e_{j}$ and construct queries asking whether $f(e_{i})>f(e_{j})$. For WD, Taste and Rocks, we sample 500 such pairs for each of the features. Since the TG dataset has a total of 70 features, we limit the test set to 100 pairs per feature. For this analysis, we have split the WD1 dataset into two parts: 80% of the entities, for each feature, are used for training the models. The remaining 20% are used as a test set. All models are fine-tuned on the training split of WD1 (apart from the baselines, which are evaluated zero-shot). This allows us to see how well the models perform on the features they were trained on (for WD1-test) and how they generalise to unseen properties.
The aim of the analysis in Table 1 is to assess whether models can be successfully fine-tuned using a relatively small training set (i.e. WD1-train), involving only well-defined numerical features. In particular, we want to test whether models which are fine-tuned on such features would also generalise to more subjective and less readily available ones, similar to the easy-to-hard generalisation that has been observed for LLMs in other tasks Hase et al. (2024). The results show that this is indeed the case, at least to some extent. Overall, we can see that Mistral-7B achieves the best results among the pointwise models, while Llama3-8B achieves the best results among the pairwise models. The pointwise Mistral-7B model outperforms the pairwise Mistral-7B model, which is surprising given that pairwise models generally perform better in ranking tasks. The performance of the models across different features is not always consistent. Each model achieves close to random chance on some of the features, but the features where one model performs poorly are not always the same features where other models perform poorly. However, for bitterness and roughness, the accuracy of all models is below 61%. Furthermore, sourness and organisation also stand out as being more challenging. Regarding the baselines, GPT-4 generally performs better than the fine-tuned models. Nonetheless, the pairwise Llama3-8B and pointwise Mistral-7B models outperform GPT-4 in several cases. ChatGPT performs worse on most features, but achieves the best results for books.
@@
The fine-tuned models in Table 1 generally underperform GPT-4. This may be partially explained by the fact that a small training set was used, which moreover only covered numerical features. In Table 2, we evaluate the impact of using different training sets. For this analysis, we use the pairwise Llama3-8B model, the best fine-tuned model in Table 1. Our focus is on seeing whether models trained on one domain can generalise to other domains. To put the results in context, we also compare with a model that is not fine-tuned but includes three in-context demonstrations (shown as few-shot); see Appendix A for details. The results are again evaluated in terms of accuracy, using the same pairwise judgements as for Table 1. WD refers to the full dataset (including both the training and test splits of WD-1).
@@
We can see in Table 2 that training on larger datasets indeed leads to considerably better results. While this is not unexpected, we can also make more striking observations. For instance, the model that was trained on WD+TG+Taste achieves strong results on Rocks, despite covering very different features. Similarly, the model that was trained on TG alone achieves strong results for both Taste and Rocks. This suggests that the fine-tuned models are indeed capable of generalising to unseen domains. However, to achieve strong results, it appears to be important that the training data contains subjective or perceptual features. In particular, training on WD alone leads to clearly worse results on Taste and Rocks. The best results in Table 2 generally outperform the GPT-4 results from Table 1. As the training and test sets cover disjoint domains, the results reflect the knowledge that is captured by the LLMs themselves, rather than knowledge that was injected during the fine-tuning process. This suggests that pre-trained LLMs capture more perceptual knowledge than it may initially appear.
Feature Ranking
Is $X$ more suitable as a present for a 10-year-old girl than $Y$? 1. lego, 2. bicycle, 3. lipgloss, 4. puppy, 5. airpods, 6. guitar, 7. helicopter, 8. Famous Five Collection - Enid Blyton, 9. broom, 10. watch, 11. gold fish, 12. salmon fish, 13. knife, 14. crocodile, 15. coffee mug, 16. Sidney Sheldon Series
Does $X$ require more urgent medical treatment than $Y$? 1. cardiac arrest, 2. brain haemmorage, 3. stroke, 4. infection, 5. flu, 6. cataract surgery, 7. cough and cold, 8. dental fillings, 9. fist bump, 10. paper cut
Is $X$ more important than $Y$ post marathon? 1. run marathon, 2. hydrate, 3. indulge in alcohol, 4. partying, 5. shopping, 6. rest, 7. skip meal, 8. snacking, 9. haircut
Is $X$ cheaper to live than $Y$? 1. Mexico City, 2. Budapest, 3. Bali, 4. Bangalore, 5. Bangkok, 6. Paris, 7. Singapore, 8. New York, 9. Los angeles, 10. Tokyo
Does $X$ burn more calories than $Y$? 1. running, 2. swimming, 3. stair climbing, 4. eating, 5. strolling, 6. tai chi, 7. pilates, 8. drinking beer, 9. sleeping, 10. watching netflix
Does the salary of $X$ exceed that of $Y$? 1. entrepreneur, 2. doctor, 3. lawyer, 4. consultant, 5. engineer, 6. scientist, 7. lecturer, 8. soldier, 9. mechanic, 10. dance teacher
Is $X$ more suitable for long-distance travel with a family than $Y$? 1. spaceship, 2. minivan, 3. aeroplane, 4. freight train, 5. local train, 6. speed boat, 7. submarine, 8. taxi, 9. quad bike, 10. motor cycle
Is $X$ more likely than $Y$ to be the first thing you do in the morning? 1. hygiene routine, 2. hitting disco, 3. watch movies, 4. making bed, 5. socialize, 6. getting ready for shopping, 7. attend in-person class, 8. have snack, 9. checking emails, 10. karaoke
Is $X$ more important than $Y$ when selecting a primary school for children? 1. alumni network, 2. parent’s income, 3. academic reputation, 4. nearby shopping centre, 5. swimming pool facilities, 6. breakfast club availability, 7. playground, 8. classroom size, 9. parking facilities, 10. school location
Is $X$ more urgent than $Y$ when considering home improvement projects? 1. bathroom renovation, 2. repairing broken windows, 3. fixing a leak, 4. installing new locks, 5. landscaping the backyard, 6. repairing door lock, 7. installing new furniture, 8. installing new light, 9. painting ceilings, 10. changing carpet
Table 5: Examples of entity rankings for selected (commonsense) features. Results were obtained with the pairwise Llama3-8B model trained on WD+TG+Rocks.
@@
Table 3 compares different strategies for generating rankings. The pointwise model can be directly used for this purpose. For the pairwise model, we show results with the SVM strategy and the Count strategy. We furthermore vary the number of pairwise judgments per entity (5 or 30) for Count and SVM. For this experiment, we use the pointwise and pairwise Llama3-8B models that were trained on WD+TG+Rocks. We evaluate the different models by comparing the predicted rankings with the ground truth in terms of Spearman $\rho$. We can see that the pairwise approaches consistently outperform the pointwise model. The SVM method generally performs better than the Count method, especially in the case where 30 judgments per entity are obtained. We also compare with the GPT-3 results reported by Chatterjee et al. (2023), finding that the pairwise Llama model consistently performs best.